# Generalized and Efficient 2D Gaussian Splatting for Arbitrary-scale Super-Resolution

### v3

@@ 1 Introduction
Image Super-Resolution (SR) has been a challenging inverse problem in computer vision for many years, and a variety of deep learning-based (SR) methods have been developed in the past decade [46, 58, 31, 5, 43, 47, 56, 50, 51, 55, 6]. However, most of the existing SR models [63, 64, 45, 30, 61, 9, 15, 29, 59] are designed to operate only under integer or fixed scaling factors, i.e., ×2,×3,×4$\times$ 2$\times$ 3$\times$ 4× 2 , × 3 , × 4, due to the fact that they perform resolution upsampling by processing discrete pixels with de-convolutional layers [15] or pixel-shuffle layers [40]. As a result, it is hard for them to perform SR with decimal or even arbitrary scaling factors. Although the existing SR models could achieve arbitrary-scale zooming via adopting bicubic or bilinear interpolation, they would inevitably scarify image quality, leading to unsatisfied super-resolved results.
@@
Recently, Gaussian Splatting (GS) [22] has shown its advantages over INR-based methods, such as Neural Radiance Field (NeRF) [37], in terms of both accuracy and complexity, thanks to its high computational efficiency and powerful explicit representation capabilities. GS has been successfully applied to tasks such as 3D reconstruction [17, 4], representation [38, 16], editing [42, 11], generation [33, 12], etc. While some studies have explored the use of GS for low-level vision tasks, e.g., image deblurring [8, 65], image compression [62], video super-resolution [53], these models overfit on the repetitive patterns from multi-frames in a video clip or multi-view cameras, or overfit on a single image to complete the image compression task. None of them could be applied to the image SR task, since overfitting a model on a single image cannot meet the generalization requirements of SR, where a single model should be used to super-resolve any input low-resolution (LR) image to its high-resolution (HR) counterpart.
Instead of overfitting the model to a fixed scenario, we switch from the optimization-based manner of original 3D GS [22] to a learning-based method and present GSASR (Gaussian Splatting based Arbitrary-scale Super-Resolution). Specifically, we first obtain the deep image features by encoding the input LR image with a general SR backbone [32, 63]. Secondly, inspired by DETR [3], we introduce learnable Gaussian embedding to output 2D Gaussians conditioned on the extracted image features. To enhance interaction among Gaussian embeddings, we perform window self-attention [34] so that they can adapt to each other. Then, the Gaussian embeddings are projected to predict the corresponding properties of 2D Gaussians, including the opacity, position, standard deviation, correlation coefficient, and peak RGB values. These 2D Gaussians depict the image in the continuous Gaussian space. To render 2D Gaussians to obtain the super-resolved image with arbitrary scale, we implement an efficient differentiable GPU/CUDA-based rasterization, which receives the predicted 2D Gaussians with an upsampling scale vector as condition. The scale vector controls the sampling density, and all Gaussians are processed efficiently in parallel.
@@ 2 Related Work
Very recently, GaussianSR [19] was proposed to employ GS for the ASR task. However, it uses only utilizes the pixel positions of LR image to substitute the center positions of Gaussians, heavily limiting the representation capability of Gaussians. In contrast, our proposed GSASR predicts all of the parameters of Gaussians, and we present an efficient differentiable 2D GPU/CUDA-based scare-aware rasterization to largely reduce the computational cost, leading to an effective and efficient model for accomplishing the ASR task.
@@ 3 Method
In this section, we first describe the concept and formulation of 2D Gaussian Spatting (GS), then present the model architecture and training details of our GSASR method.
@@
2D GS Representation. Our key insight is to represent an image using contiguous Gaussians, which can fit various frequency components of an image due to the powerful expressive properties of GS. Specifically, considering an LR image $I_{LR}$ of size $H\times W$, we represent it using $N$ 2D Gaussians, denoted by $\mathcal{G}=\{G_{1},G_{2},...,G_{N}\}$. Each Gaussian maintains properties of opacity $\alpha~{}\in[0,1]^{1}$, center position $\mu=\{\mu_{x}~{}\in\mathbb{R}^{1},\mu_{y}~{}\in\mathbb{R}^{1}\}$, standard deviation $\sigma=\{\sigma_{x}~{}\in(0,\infty)^{1},\sigma_{y}~{}\in(0,\infty)^{1}\}$, correlation coefficient $\rho~{}\in[-1,1]^{1}$, and normalized peak RGB value $c~{}\in[0,1]^{3}$. We define the RGB color at position $(x,y)$ contributed by Gaussian $G$ as $G(x,y)$ (here we omit the index of Gaussians for simplicity of expression):
@@
where $\Delta x=x-\mu_{x}$ and $\Delta y=y-\mu_{y}$ decide the position of Gaussians on an image, and the standard deviation $\sigma_{x}$, $\sigma_{y}$ and the correlation coefficient $\rho$ indicate the shape and direction of Gaussian. The pixel color of the LR image at position $(x\in[0,W-1],y\in[0,H-1])$ is represented as the summarization of all $N$ Gaussians:
@@ 6       For each pixel $(x,y)$ in $I_{SR}$ do
Gaussian functions $G$ are infinite and continuous on the 2D space, enabling discretely sampling at arbitrary positions. With a given scaling factor $s$, to obtain an SR image $I_{SR}$ of size $sH\times sW$, we sample the color at position $(x\in[0,sW-1],y\in[0,sH-1])$ with the following scale-aware rasterization:
$I_{SR}(x,y;s)=\sum_{i=1}^{N}G_{i}(\frac{x}{s},\frac{y}{s}).\vspace{-0.8em}$ (4)
Table 1: Quantitative comparison between representative ASR models and our GSASR. All models use the same EDSR-backbone [32] as the feature extraction encoder and are tested on DIV2K and LSDIR [28] datasets [41] with scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$, ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$. The best results are highlighted in red. The PSNR and SSIM indices are computed in the Y channel of Ycbcr space.
@@ 3.2 GSASR: Architecture and Modules
Gaussian Embedding. In the GS decoder, we aim to represent the LR image through $N$ Gaussian embeddings $\mathbf{E}\in\mathbb{R}^{N\times d}$ conditioned on $\mathbf{F}$, where $d$ is the embedding dimension. As for the number of Gaussian embeddings $N$, a straightforward idea is to set it to a fixed number so that the $N$ independent learnable Gaussian embeddings interact with image features to output 2D Gaussians $\mathcal{G}$. However, this approach suffers from low effectiveness and computational efficiency. As the LR images exhibit variable sizes in the test phase, rendering a fixed number of 2D Gaussians will either waste computational resources when handling LR image with very small size, or lead to deficient representation when processing larger LR images. Another possible strategy is to set $N\propto s$ under different scaling factord $s$. However, it requires the Gaussian decoder to handle varying sizes of Gaussian embeddings, while the size of LR features is fixed in the training stage. In addition, as the scaling factor increases, more Gaussians are introduced, leading to longer inference time and higher GPU memory cost.
@@
where $Q$ is from the Gaussian embeddings $\mathbf{E}$, $K$ and $V$ are from image features $\mathbf{F}$. The $i$-th embedding $\mathbf{E}_{i}$ at position $p_{i}$ interacts with the $j$-th pixel feature within the window. $B$ is inherited from the Swin Transformer [34], serving as a learnable position encoding to indicate the relative position between Gaussian embeddings and LR features.
Gaussian Interaction Block. In the condition injection block, each Gaussian embedding $\mathbf{E}_{i}$ is processed independently, resulting in a lack of information exchange among different Gaussian embeddings. To address this issue, we stack $L$ Gaussian interaction blocks to enhance the interaction among Gaussian embeddings. The Gaussian interaction block is designed based on the Swin Transformer architecture [34], which iteratively stacks the window attention layers and shifted window attention layers. In attention layers, we perform self-attention for embeddings whose reference positions are located in the same window, similar to the condition injection block. Besides, we introduce a cross-attention layer to make embeddings scale-aware so that the Gaussian embeddings can learn to adjust the properties according to the given scale factor.
@@ 4.1 Experimental Setup
Implementation Details. We utilize the widely-used DIV2K [41] dataset as training set. We follow [10, 2] to fix LR patches to $48\times 48$. The scaling factor $s$ is randomly selected within the range $[1.0,4.0]$. We first crop $48s\times 48s$ GT patch from the original full-size GT image, and then apply bicubic down-sampling [2] to obtain the corresponding LR patches. Following LIIF [10], we employ EDSR [32] and RDN [64] as the image encoder backbones to extract features from the input LR images. We train our GSARS on $4$ NVIDIA A100 GPUs for $500,000$ iterations with batch size 64. The initial learning rate is $2e^{-4}$, and halves at $250,000$, $400,000$, $450,000$, $475,000$ iterations. The Adam [25] optimizer is utilized. To speed up the convergence of the training progress, we set the number of warm-up iterations to $2,000$. The window size $k$ in Condition injection block and Gaussian Interaction Block is set to $12$, we set the proportion parameter $m=16$, the dimension of Gaussian embedding $d=180$, and stack $L=6$ Gaussian interaction blocks.
@@
Evaluation Protocols. We evaluate the competing models on Set5 [1], Set14 [57], DIV2K100 [41], Urban100 [21], BSDS100 [35], Manga109 [36], General100 [15], and LSDIR [28] datasets under scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$. To validate the generalization on higher scaling factors, we further report the results with ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$ scaling factors on DIV2K100 [41] and LSDIR [28] datasets. The PSNR, SSIM [48], LPIPS [60] and DISTS [13] indices are used for performance evaluation. Note that some existing methods [10, 2] calculate PSNR on RGB channels for DIV2K [41], but on the Y channel of YCbCr space for other datasets. To unify the testing protocols, we calculate PSNR/SSIM on the Y channel across all datasets and all competing methods for comparison. For a fair comparison, we download all competing models from their official websites, utilize the same data to generate SR results and the same evaluation codes to compute the metrics.
@@ 4.2 Experiment Results
Ablation Study. We conduct ablation studies on (1) the number of Gaussians $N$, (2) the functionality of the reference position, (3) the rasterization ratio $r$, (4) the dimension $d$ of Gaussian embedding, abd (5) the window size $k$ in condition injection block and Gaussian Interaction block. Details are presented in the supplementary file.

### v4

@@ 1 Introduction
Image Super-Resolution (SR) has been a challenging inverse problem in computer vision for many years, and a variety of deep learning-based (SR) methods have been developed in the past decade [46, 58, 31, 5, 43, 47, 56, 50, 51, 55, 6]. However, most of the existing SR models [63, 64, 45, 30, 61, 9, 15, 29, 59] are designed to operate only under integer or fixed scaling factors, i.e., ×2,×3,×4$\times$ 2$\times$ 3$\times$ 4× 2 , × 3 , × 4, due to the fact that they perform resolution upsampling by processing discrete pixels with de-convolutional layers [15] or pixel-shuffle layers [40]. As a result, it is hard for them to perform SR with decimal or even arbitrary scaling factors. Although the existing SR models could achieve arbitrary-scale zooming via adopting bicubic or bilinear interpolation operations, they would inevitably lead to unsatisfied or over-smoothing super-resolved results.
@@
Recently, Gaussian Splatting (GS) [22] has shown its advantages over INR-based methods, such as Neural Radiance Field (NeRF) [37], in terms of both accuracy and complexity, thanks to its high computational efficiency and powerful explicit representation capabilities. GS has been successfully applied to tasks such as 3D reconstruction [17, 4], representation [38, 16], editing [42, 11], generation [33, 12], etc. While some studies have explored the use of GS for low-level vision tasks, e.g., image deblurring [8, 65], image compression [62], video super-resolution [53], these models overfit on the repetitive patterns from multi-frames in a video clip or multi-view cameras, or overfit on a single image to complete the image compression task. None of them could be applied to the image ASR task, since overfitting a model on a single image cannot meet the generalization requirements of ASR, where a single model should be used to super-resolve any input low-resolution (LR) image with any scales to its high-resolution (HR) counterpart.
Instead of overfitting the model to a fixed scenario, we switch from the optimization-based manner of original 3D GS [22] to a learning-based method and present GSASR (Gaussian Splatting based Arbitrary-scale Super-Resolution). Specifically, we first obtain the deep image features by encoding the input LR image with a general SR backbone [32, 63]. Secondly, inspired by DETR [3], we introduce learnable Gaussian embeddings to output 2D Gaussians conditioned on the extracted image features. To enhance interaction among Gaussian embeddings, we perform window self-attention [34] so that they can adapt to each other. Then, the Gaussian embeddings are projected to predict the corresponding properties of 2D Gaussians, including the opacity, position, standard deviation, correlation coefficient, and peak RGB values. These 2D Gaussians depict the image in the continuous Gaussian space. To render 2D Gaussians to obtain the super-resolved image with arbitrary scale, we implement an efficient differentiable GPU/CUDA-based rasterization, which receives the predicted 2D Gaussians with an upsampling scale vector as condition. The scale vector controls the sampling density, and all Gaussians are processed efficiently in parallel.
@@ 2 Related Work
Very recently, GaussianSR [19] was proposed to employ GS for the ASR task. However, it only utilizes the pixel positions of LR image to substitute the center positions of Gaussians, heavily limiting the representation capability of Gaussians. In contrast, our proposed GSASR predicts all of the parameters of Gaussians, and we present an efficient differentiable 2D GPU/CUDA-based scare-aware rasterization to largely reduce the computational cost, leading to an effective and efficient model for accomplishing the ASR task.
@@ 3 Method
In this section, we first describe the concept and formulation of 2D Gaussian Splatting (GS), then present the model architecture and training details of our GSASR method.
@@
2D GS Representation. Our key insight is to represent an image using continuous Gaussians, which can fit various frequency components of an image due to the powerful expressive properties of GS. Specifically, considering an LR image $I_{LR}$ of size $H\times W$, we represent it using $N$ 2D Gaussians, denoted by $\mathcal{G}=\{G_{1},G_{2},...,G_{N}\}$. Each Gaussian maintains properties of opacity $\alpha~{}\in[0,1]$, center position $\mu=\{\mu_{x}~{}\in\mathbb{R},\mu_{y}~{}\in\mathbb{R}\}$, standard deviation $\sigma=\{\sigma_{x}~{}\in(0,\infty),\sigma_{y}~{}\in(0,\infty)\}$, correlation coefficient $\rho~{}\in[-1,1]$, and normalized peak RGB value $c~{}\in[0,1]$. We define the RGB color at position $(x,y)$ contributed by Gaussian $G$ as $G(x,y)$ (here we omit the index of Gaussians for simplicity of expression):
@@
where $\Delta x=x-\mu_{x}$ and $\Delta y=y-\mu_{y}$ decide the position of Gaussians on an image, and the standard deviation $\sigma_{x}$, $\sigma_{y}$ and the correlation coefficient $\rho$ indicate the shape and direction of Gaussian. The pixel color of the LR image at position $(x\in[0,W-1],y\in[0,H-1])$ is represented as the summation of all $N$ Gaussians $\mathcal{G}$:
@@
Each Gaussian function $G_{i}$ is infinite and continuous on the 2D space, enabling discretely sampling at arbitrary positions. With a given scaling factor $s$, to obtain an SR image $I_{SR}$ of size $sH\times sW$, we sample the color at position $(x\in[0,sW-1],y\in[0,sH-1])$ with the following scale-aware rasterization:
$I_{SR}(x,y;s)=\sum_{i=1}^{N}G_{i}(\frac{x}{s},\frac{y}{s}).\vspace{-0.8em}$ (4)
@@ 6       For each pixel $(x,y)$ in $I_{SR}$ do
Table 1: Quantitative comparison between representative ASR models and our GSASR. All models use the same EDSR-backbone [32] as the feature extraction encoder and are tested on DIV2K and LSDIR [28] datasets [41] with scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$, ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$. The best results are highlighted in red. PSNR/SSIM metrics are computed on the Y channel of Ycbcr space.
@@ 3.2 GSASR: Architecture and Modules
Gaussian Embedding. In Gaussian decoder, we aim to represent the LR image through $N$ Gaussian embeddings $\mathbf{E}\in\mathbb{R}^{N\times d}$ conditioned on $\mathbf{F}$, where $d$ is the embedding dimension. As for the number of Gaussian embeddings $N$, a straightforward idea is to set it to a fixed number so that the $N$ independent learnable Gaussian embeddings interact with image features to output Gaussians $\mathcal{G}$. However, this approach will either be ineffective or lead to low computational efficiency. As the LR images exhibit variable sizes in the test phase, rendering a fixed number of Gaussians will either waste computational resources when handling LR image with very small size, or lead to deficient representation when processing larger LR images. Another possible strategy is to set $N\propto s$ under different scaling factord $s$. However, it requires the Gaussian decoder to handle varying sizes of Gaussian embeddings, while the size of LR features is fixed in the training stage. In addition, as the scaling factor increases, more Gaussians are introduced, leading to longer inference time and higher GPU memory cost.
@@
where $Q$ is from the Gaussian embeddings $\mathbf{E}$, $K$ and $V$ are from image features $\mathbf{F}$. The $i$-th embedding $\mathbf{E}_{i}$ at position $p_{i}$ interacts with the $j$-th pixel feature within the window. $B$ is inherited from Swin Transformer [34], serving as a learnable position encoding to indicate the relative position between Gaussian embeddings and LR features. The window size of window-based cross attention is set to the same $k$ value as that of the partitioned window LR features.
Gaussian Interaction Block. In the condition injection block, each Gaussian embedding $\mathbf{E}_{i}$ is processed independently, resulting in a lack of information exchange among different Gaussian embeddings. To address this issue, we stack $L$ Gaussian interaction blocks to enhance the interaction among Gaussian embeddings. The Gaussian interaction block is designed based on the Swin Transformer architecture [34], which iteratively stacks the window attention layers and shifted window attention layers. In attention layers, we perform self-attention for embeddings whose reference positions are located in the same window, similar to the condition injection block. Besides, we introduce a cross-attention layer to make embeddings scale-aware so that the Gaussian embeddings can learn to adjust the properties according to the given scale factor. The window size of window attention layer is set to the same $k$ value as that of the partitioned window LR features.
@@ 4.1 Experimental Setup
Implementation Details. We utilize the widely-used DIV2K [41] dataset as training set. We follow [10, 2] to fix LR patches to $48\times 48$. The scaling factor $s$ is randomly selected within the range $[1.0,4.0]$. We first crop $48s\times 48s$ GT patch from the original full-size GT image, and then apply bicubic down-sampling [2] to obtain the corresponding LR patches. Following LIIF [10], we employ EDSR [32] and RDN [64] as the image encoder backbones to extract features from the LR images. We train GSASR on $4$ NVIDIA A100 GPUs for $500,000$ iterations with batch size 64. The initial learning rate is $2e^{-4}$, and halves at $250,000$, $400,000$, $450,000$, $475,000$ iterations. The Adam [25] optimizer is utilized. To speed up the convergence of the training progress, we set the number of warm-up iterations to $2,000$. The window size $k$ in window LR features, condition injection block and Gaussian interaction block is set to $12$, the proportion parameter $m$ in Gaussian embedding is set to $16$, the dimension $d$ of Gaussian embedding is set to $180$, and we stack $L=6$ Gaussian interaction blocks.
@@
Evaluation Protocols. We evaluate the competing models on Set5 [1], Set14 [57], DIV2K100 [41], Urban100 [21], BSDS100 [35], Manga109 [36], General100 [15], and LSDIR [28] datasets under scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$. To validate the generalization on higher scaling factors, we further report the results with ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$ scaling factors on DIV2K100 [41] and LSDIR [28] datasets. The PSNR, SSIM [48], LPIPS [60] and DISTS [13] metrics are used for performance evaluation. Note that some existing methods [10, 2] calculate PSNR on RGB channels for DIV2K [41], but on Y channel of YCbCr space for other datasets. To unify the testing protocols, we calculate PSNR/SSIM on the Y channel across all datasets and all competing methods for comparison. For a fair comparison, we download all competing models from their official websites, utilize the same data to generate SR results and the same evaluation codes to compute the metrics.
@@ 4.2 Experiment Results
Ablation Study. We conduct ablation studies on (1) the number of Gaussians $N$, (2) the functionality of the reference position, (3) the rasterization ratio $r$, (4) the dimension $d$ of Gaussian embedding, abd (5) the window size $k$ in partitioned window LR features, condition injection block and Gaussian interaction block. Details are presented in the supplementary file.